# VALUE: Value-Aware Large Language Model for Query Rewriting via Weighted Trie in Sponsored Search

### v1

@@ 4.5. Exploration of Hyperparameter Configurations
In this section, we provide a concise analysis of the impact of the hyperparameter $\theta$ on system performance. As shown in Table 2, an increase in $\theta$ results in a decrease in relevance. Figure 3 visually represents the dynamics of this trade-off. Initially, optimizing for value leads to an increase in RPM; however, when relevance becomes significantly low, RPM ultimately decreases. This underscores the critical importance of meticulous hyperparameter tuning to ensure alignment with specific application objectives. For a detailed exploration of the hyperparameter $\theta$, please refer to Appendix A.
@@ 4.8. Case Study
Figure 4. Case Study of VALUE Framework
[…]
In this section, as illustrated in Figure 4, we examine a specific case involving the user query ”fatty’s drink.” Previously, without the use of LLMs, our system could only match the query at the keyword level, specifically ”drink.” With the integration of LLMs, we gained a deeper understanding of the user’s intent, identifying a preference for low-sugar, low-calorie beverages. Additionally, by employing a weighted trie, we ensured that the advertisements presented were not only closely aligned with the user’s intent but also of high value.
@@ 5. Conclusion
In this study, we introduced VALUE, an innovative framework designed to enhance query rewriting in sponsored search advertising by integrating value-awareness into the generation process. Our approach addresses the significant challenge of producing rewrites that are not only semantically relevant but also economically valuable—a gap that existing methods often fail to bridge. By embedding value information directly into the LLM’s decoding process through a novel weighted trie structure, we effectively steer generation towards high-value bidwords without substantially compromising relevance.
Comprehensive experiments underscore the effectiveness of VALUE in enhancing the economic attributes of bidword generation. Offline evaluations demonstrate a significant increase in eCPM, outperforming traditional methods while preserving high relevance scores. The improvement in Spearman rank correlation confirms the superior ranking quality achieved through our fine-grained value alignment. Furthermore, online A/B testing validates the real-world efficacy of VALUE, highlighting notable improvements in key performance indicators such as cost and RPM. Our framework’s versatility also suggests potential applications beyond sponsored search advertising, wherever the generation space is fixed and reward signals are present.
Appendix A Hyperparameter Details
In this appendix, we delve into the specifics of the hyperparameter $\theta_{d}$, which is designed to vary with the depth of the trie. The selection of $\theta_{d}$ is critical in modulating the influence of value relative to relevance as the trie depth changes. The configurations explored in our research are as follows:
  • $\theta_{1}$: [0, 0, 0, 0, 0, 0, 0, …], indicating that the weighted trie is ineffective and the original LLM reasoning result is used.
  • $\theta_{2}$: [1, 1, 1, 1, 1, 1, 1, …], implying equal weighting of relevance and value at each depth level.
  • $\theta_{3}$: [1, 2, 3, 4, 5, 6, 7, …], where the weight increases linearly with depth, emphasizing value as the trie deepens.
  • $\theta_{4}$: [1, 2, 4, 8, 16, 32, 64, …], representing an exponential increase in weight with depth, further prioritizing value.
  • $\theta_{5}$: [2, 4, 8, 16, 32, 64, 128, …], offering a more pronounced exponential tendency compared to $\theta_{4}$.
These configurations allow for nuanced control over the relevance-value trade-off, enabling the system to adapt its focus based on the hierarchical structure of the weighted trie. The choice of $\theta_{d}$ thus plays a pivotal role in optimizing the overall performance of the inference system.

### v2

@@ 4.5. Exploration of Hyperparameter Configurations
In this section, we conduct a thorough investigation into the impact of varying the hyperparameter $\Theta$ on system performance. This parameter, which scales with the depth of the Weighted Trie, modulates the trade-off between relevance and commercial value at different stages of the inference process. Specifically, we evaluate five distinct configurations (see Table 2):
  • $\Theta_{1}$: Assigns a value weight of 0 throughout, effectively disabling the Weighted Trie and relying solely on the original LLM inference output.
  • $\Theta_{2}$: Assigns equal weighting to relevance and value at every depth of the Weighted Trie.
  • $\Theta_{3}$: Increases the value weight linearly as a function of Weighted Trie depth.
  • $\Theta_{4}$: Increases the value weight exponentially with Weighted Trie depth, significantly prioritizing value in the deeper inference stages.
  • $\Theta_{5}$: Further increases the baseline value weights at all depths established by $\Theta_{4}$.
Table 2 and Figure 3 present the performance metrics under different $\Theta$ settings across both online and offline evaluations. Key observations reveal that increasing the value weight consistently compromises relevance. Critically, however, excessive value weighting (as in $\Theta_{4}$ and $\Theta_{5}$) not only severely degrades relevance but also diminishes RPM and hitrate. Within the sponsored search context, the objective of query rewriting is to maximize the value of rewrites while satisfying a minimum relevance criterion. The $\Theta_{3}$ configuration addresses this by directing VALUE to prioritize highly relevant bidwords during the initial token generation stages, subsequently shifting focus to high-value bidwords later in the process. This strategy strikes an optimal balance between relevance and value by first ensuring the generation of a more relevant prefix. Our findings underscore that meticulous hyperparameter tuning is paramount for aligning model behavior with specific sponsored search objectives.
@@ 4.8. Case Study
Figure 4. Case Study of VALUE Framework:
In this section, as illustrated in Figure 4, we examine a specific case involving the user query ”drink for fat person”. Previously, without the use of LLMs, our system could only match the query at the keyword level, specifically ”drink” . With the integration of LLMs, we gained a deeper understanding of the user’s intent, identifying a preference for low-sugar, low-calorie beverages. Additionally, by employing a Weighted Trie, we ensured that the advertisements presented were not only closely aligned with the user’s intent but also of high commercial value.
@@ 5. Conclusion
This study introduces VALUE, an innovative framework that addresses the critical challenge of producing rewrites that simultaneously optimize semantic relevance and commercial value, which is a critical limitation pervasive in existing methods. VALUE incorporates real-time value signals into LLM decoding through a novel Weighted Trie and applies WDPO-based alignment that prioritizes bidword pairs with large value gaps during training. These components jointly guide generation toward high-value rewrites without compromising semantic relevance.
Extensive experiments confirm the framework’s effectiveness. Offline tests yield substantial eCPM gains while preserving relevance, and online A/B experiments report significant improvements in cost and Revenue Per Mille. The modular design of VALUE also supports other multi-reward tasks, for example video recommendation in which click-through rate and watch-time completion must be balanced within a fixed generation space and dynamically weighted objectives.